Multiple-Level Break-Even Analysis
Nielsen Associates provides marketing services for a number of small manufacturing firms. Nielsen receives a commission of 10 percent of sales. Operating costs are as follows:

Unit-level costs $0.02 per sales dollar
Sales-level costs $200 per sales order
Customer-level costs $800 per customer per year
Facility-level costs $60,000 per year

(a) Determine the minimum order size in sales dollars for Nielsen to break even on an order.
$Answer



(b) Assuming an average customer places five orders per year, determine the minimum annual sales required to break even on a customer.
$Answer

(c) What is the average order size in (b)?
$Answer

(d) Assuming Nielsen currently serves 100 customers, with each placing an average of five orders per year, determine the minimum annual sales required to break even.
$Answer

(e) What is the average order size in (d)?
$Answer

(f) Explain the differences in the answers to (a), (c), and (e).

In multiple customer firms the break-even point decreases as the number of customers increases.

Even if individual orders have a positive contribution, some customers may be unprofitable.

In the long-run the most important costs are facility level costs.

The most important costs to cover are unit level costs.
Answer #1

(a) Determine the minimum order size in sales dollars for Nielsen to break even on an order.

An order involves 2 following direct cost associated with it

Unit-level costs

Sales-level costs

To break even it should cover above two costs

Lets assume X is the minimum order size

X = $200 + $0.02X

0.98 X = 200

X = $200/0.98

Minimum Order Size

$                                         204.08

(b) Assuming an average customer places five orders per year, determine the minimum annual sales required to break even on a customer

Costs associated with a customer involves following costs

Unit-level costs

$0.02 per sales dollar

Sales-level costs

$200 per sales order

Customer-level costs

$800 per customer per year

Lets assume that X is the minimum annual sales to a customer

X = 800 + 5*200 + 0.02*X

0.98X = 1800

X = 1800/0.98

$                                     1,836.73

Minimum annual sales required to break even on a customer

$                                     1,836.73

(c) What is the average order size in (b)?

Average order size in (b) = Minimum annual sales required to break even on a customer/ Average orders placed by a customer

=1836.73/5

$                                                         367.35

(d) Assuming Nielsen currently serves 100 customers, with each placing an average of five orders per year, determine the minimum annual sales required to break even.

Costs associated with a customer involves following costs

Unit-level costs

$0.02 per sales dollar

Sales-level costs

$200 per sales order

Customer-level costs

$800 per customer per year

Facility-level costs

$60,000 per year

Lets assume Y is Minimum annual sales required to break even

Y = 60,000 + 800*100 + 200*100*5 + 0.02*Y

0.98Y = 240000

Y = 240000/0.98

Minimum annual sales required to break even

$                                 244,897.96

(e) What is the average order size in (d)?

Average Order size = Average Annual Sales/Number of orders

=244,897.96/(100*5) = $48,979.6

f. The difference in answers is due to multiple level break even like cost centre accounting and the total cumulative cost is in last part.